OpenStudy (anonymous):

The table shows the American League batting champions for the years 1999–2003. The sum of these five batting averages is 1.754. Write and solve an equation to determine the batting average x for Bill Mueller, the 2003 batting champion.

OpenStudy (anonymous):

American League Leading Hitters (1992–1988) Year Player Team Batting Average 1999 Nomar Garciaparra Boston .357 2000 Nomar Garciaparra Boston .372 2001 Ichiro Suzuki Seattle .350 2002 Manny Ramirez Boston .349 2003 Bill Mueller Boston x A. .357 + .372 + .350 + .349 + x = 1.754; .326 B. .357 + .372 + .350 + .349 + x = 1.754; .316 C. x + 1.754 = .357 + .372 + .350 + .349; .326 D. 1.718 – x = .357 + .372 + .350 + .349; .330